OpenStudy (anonymous):

Find the zeros of the polynomial function. f(x) = x^3 + 5x^2 - 4x - 20

OpenStudy (kinggeorge):

Do you know how to factor by grouping?

OpenStudy (anonymous):

I do but when three is x^3 it confuses me

OpenStudy (kinggeorge):

I'll do the factor by grouping part for you. \[x^3+5x^2-4x-20=x^2(x+5)-4(x+5)=(x^2-4)(x+5)\]Can you factor \(x^2-4\)?

OpenStudy (anonymous):

(x - 2)(x +2)

OpenStudy (kinggeorge):

Perfect. That means your polynomial factors like \[(x-2)(x+2)(x+5)\]What does this tell you about the zeros?

OpenStudy (anonymous):

so they would be -2,2,5

OpenStudy (kinggeorge):

Almost. Your zeros would be -2,2,-5

OpenStudy (anonymous):

so I have to do this first x+5=0 x=-5

OpenStudy (kinggeorge):

Right.

OpenStudy (anonymous):

then I get my zeros

OpenStudy (anonymous):

thank you I just have a hard time factoring when the equation is like that

OpenStudy (kinggeorge):

You're welcome.